Good day. I'd be thankful if I can ask anyone to help me on this problem, since the answers that I'm getting barely make sense even to me. It goes like this:
A fruit dealer placed together $1.50-a-dozen apples and $1.20-a-dozen apples, putting 2 of the dealer kind for every 3 of the cheaper kind, and then sold them at $1.32-a-dozen. How many apples of the dealer kind did the vendor dispose if he sold $66.00 worth of apples?
Once again, any help would be greatly appreciated. Thanks! :)

Answer by mananth(16949)

Let the number of dealer apples he sold be x
then the number of dozens of aplles of cheaper variety will be 3x/2 ( ratio is 2:3)
the price is $1.5 /dozen. So he gets 1.5x from these apples
The price for cheaper quality is $1.2 . So he gets 1.2 *3x/2 from these apples
1.5x+1.2 * 3x/2 = 66
1.5x 3.6x/2= 66
3x+3.6x / 2 = 66
6.6= 132
x= 132/6.6
x= 20 dozens ------ dealer apples